Weather warm and sunny 26 degrees.
Hive #1 – inspected by Liz & Rowan. Queen seen and marked Yellow. (replacement for 2016 white queen)  Queen still has damaged abdomen but no supercedure cells and brood on 6 frames. Empty frame moved inside pollen barrier. Varroa = 52.
Hive #2 – Wednesday inspection.
Hive #2A – Wednesday inspection.
Hive #3 – brood box not checked as QC waiting to emerge. Wet super taken off. Next inspection 15 July. Varroa = 4.
Hive #4 – inspected by Liz & Rowan. Queen not seen. Brood on 5 frames. Dummy board removed. Varroa = 4.
Hive #5 – not inspected as waiting for Queen to emerge. Varroa count not recorded.
Hive #6 – Wednesday inspection.
Hive # 7 – inspected by Graham. Queen not seen but brood in all stages. Brood + half. Transferred four frames full of stores from half brood to super and replaced with empty drawn frames, to create more room for queen to lay. Varroa = 1.
Hive #8 – inspected by Peter. Brood + half. Queen seen. Brood on 9 frames. Super added above QX and Rhombus Board put under top super. Happy healthy colony. Varroa not recorded.